OpenStudy (anonymous):

f(x) = 10 f(x)=-5x-3

OpenStudy (anonymous):

Want medals Please help!!!

OpenStudy (blurbendy):

f(x) = -5x - 3, so substitute that into the first equation and solve for x

OpenStudy (anonymous):

I tried :/

OpenStudy (blurbendy):

-5x - 3 = 10 -5x = 13 (add 3 to both sides) what is x equal to now?

OpenStudy (blurbendy):

did you get it?

OpenStudy (anonymous):

im still putting it together

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

What about f(x) =10 F(x)=-2x+4

OpenStudy (blurbendy):

no, -5x = 13 x = -13 / 5

OpenStudy (blurbendy):

same thing -2x + 4 = 10 -2x = 6 what does x equal?

OpenStudy (anonymous):

-3?

OpenStudy (blurbendy):

yup :)

OpenStudy (anonymous):

Thankss <3

OpenStudy (blurbendy):

np, anytime!